# Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model

### 03 Mar 2026 v1

@@ Abstract.
In this paper, we study an intertemporal utility maximization problem in which an investor chooses consumption and portfolio strategies in the presence of a stochastic factor and a no-borrowing constraint. In the spirit of the Kim–Omberg model, the stochastic factor represents the excess return of the risky asset and follows an Ornstein–Uhlenbeck process, capturing the mean reversion of expected excess returns—a feature well supported by empirical evidence in financial markets. The investor seeks to maximize expected utility from consumption, subject to the constraint that wealth remains nonnegative at all times. To address the dynamic no-borrowing constraint, we use Lagrange duality to transform the primal problem into a singular control problem in the dual space. We then characterize the solution to the dual singular control problem via an auxiliary two-dimensional optimal stopping problem featuring stochastic volatility, and subsequently retrieve the primal value function as well as the optimal portfolio and consumption plans. Finally, a numerical study is conducted to derive economic and financial implications.
@@ 1. Introduction
Investment opportunities vary over time because the expected excess return of the risky asset is driven by a stochastic factor $(\beta_{t})_{t}$. Empirical evidence (see, e.g., [14] and [37]) suggests that expected excess returns are predictable and mean-reverting. Motivated by this finding, we model $(\beta_{t})_{t}$ as a mean-reverting Ornstein–Uhlenbeck process, following the so-called Kim–Omberg model introduced in [27]. Furthermore, we allow the Brownian motion driving this stochastic factor to be correlated with the Brownian motion driving the risky asset’s returns.
We derive optimal consumption and portfolio policies, as well as regularity results for the value function. As discussed in more detail below, this is achieved by means of a Lagrange duality approach, which allows us to connect the (primal) dynamic optimization problem with a no-borrowing constraint to a singular stochastic control problem. The latter is then analyzed through an auxiliary genuinely two-dimensional optimal stopping problem, in which one of the state variables exhibits stochastic volatility. In particular, we establish continuous differentiability of the optimal stopping problem’s value function and characterize the optimal stopping time in terms of an excess-return-dependent free boundary. These properties are then instrumental in obtaining the complete solution to the original optimal consumption and portfolio choice problem.
@@
Inspired by [12], [22], and the more recent [24], we instead adopt a duality-based approach to handle the no-borrowing constraint. Instead of enforcing $X_{t}\geq 0$ dynamically at every point in time, we reformulate it as a single static budget constraint. This transformation is achieved by introducing a non-increasing, càdlàg process $(D_{t})_{t}$ that acts as an endogenous dynamic Lagrange multiplier designed to ensure $X_{t}\geq 0$ at all times $t\geq 0$ almost surely. This dual formulation allows us to express the problem in terms of an auxiliary process $(Z_{t})_{t}$ rather than the wealth process itself. Consequently, the dual problem is recast as a two-dimensional singular control problem in the state variables $(Z_{t},\beta_{t})_{t}$, where $(Z_{t})_{t}=(Z_{t}^{D})_{t}$ is controlled through the monotone control $(D_{t})_{t}$. Crucially, because the stochastic factor $(\beta_{t})_{t}$ directly drives the diffusion of the dual state, the system inherently features stochastic volatility. It is worth noticing that directly approaching the dynamic programming principle (Hamilton-Jacobi-Bellmann, HJB) equation related to the primal problem is particularly challenging in our setting, because of a non-geometric wealth dynamics due to the presence of labor income. As a matter of fact, in many classical settings, the wealth process is geometric, and the positivity of wealth is automatically guaranteed. In presence of power utility functions, this geometric structure also allows the value function to be scaled with wealth, reducing the HJB equation to a one-dimensional nonlinear ODE (see, e.g., [32] and [31] for earlier works or the recent [18] and [19]). In our setting, however, the usual scaling arguments break down, the no-borrowing constraint must be enforced explicitly, and a duality approach is needed (see also Remark 2.3 below).
@@
By connecting the optimal stopping problem back to the dual singular control problem, we show that the dual value function is precisely the integral of the stopping value function with respect to the first component of the state process. This critical relationship allows us to uniquely characterize the optimal singular control $(D^{*}_{t})_{t}$, which is the minimal process (à la Skorokhod) that keeps the dual state process $(Z_{t},\beta_{t})_{t}$ within the region $\{(z,\beta):z<z^{*}(\beta)\}$. Finally, we prove that strong duality holds, demonstrating that the primal value function can be recovered from the dual value function, and vice versa. Using these results, we are then able to retrieve the optimal consumption plan $(c^{*}_{t})_{t}$, the optimal portfolio strategy $(\pi^{*}_{t})_{t}$, and the optimal wealth process $(X^{*}_{t})_{t}$, which, as required, adheres to the dynamic no-borrowing constraint.
@@
Optimal consumption and investment problems in the presence of a stochastic factor have also been widely studied. For example, [33] consider an optimal consumption problem involving stochastic interest rates, while [20] study a Merton consumption and portfolio problem with stochastic asset returns and volatilities (see also [21]). Closely related to our work is the Kim-Omberg model [27], which studies the dynamic non-myopic portfolio behavior of an investor trading a risk-free and a risky asset, with expected excess returns following a mean-reverting Ornstein-Uhlenbeck process. Extending this framework to include intermediate consumption, [40] derives exact optimal portfolio and consumption rules. However, unlike our model, [40] assumes complete markets by driving both the asset return and the stochastic factor with the same Brownian motion, does not incorporate labor income, and enforces the non-negativity constraint on wealth only at the terminal time $T$. Furthermore, [18] derive optimal consumption and investment policies in a complete market featuring a stochastic factor, modeled as a general scalar diffusion that drives investment opportunities. In their framework, the absence of labor income ensures that the wealth dynamics remain purely geometric. This in turn allows the value function to be scaled by wealth, thereby reducing the dimension of the HJB equation. Lastly, [19] consider an optimal consumption and investment model with general stochastic factors. They address optimal investment and consumption for a power utility investor using an incomplete stochastic factor model on an infinite horizon and provide a complete characterization for a finite state space. When the factor follows a diffusion process, they develop a new theoretical framework to prove existence and bound the HJB solution, verifying models like the Heston model rigorously for the first time. Conversely to our approach, [19] allows for general stochastic factors, including the Kim-Omberg setting, but assumes geometric wealth without labor income, which automatically enforces the non-negativity of wealth and crucially permits dimension reduction of the associated HJB equation.
@@ 2.1. The Financial Market and the Agent’s Problem
$dS_{t}=(r+\beta_{t})S_{t}\,dt+\sigma S_{t}\,dW_{t},\quad t>0,\quad S_{0}=s>0,$
@@
$d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
where $\kappa>0$ is the speed of mean reversion, $\overline{\beta}>0$ is the equilibrium level, and $\sigma_{\beta}>0$ is the volatility of $(\beta_{t})_{t}$. The standard Brownian motion $(W^{\beta}_{t})_{t}$ on $(\Omega,\mathcal{F},\mathbb{P})$ is correlated with $(W_{t})_{t}$ with correlation parameter $\rho\in[-1,1]$, and generates the filtration (completed by -$\mathbb{P}\text{-null sets of}\;\mathcal{F}$) $\mathbb{F}^{W^{\beta}}:=(\mathcal{F}_{t}^{W^{\beta}})_{t}$. Due to correlation, we may write, for $t\geq 0$,
$W_{t}=\rho W^{\beta}_{t}+\sqrt{1-\rho^{2}}W^{1,\perp}_{t},$
and
$W^{\beta}_{t}=\rho W_{t}+\sqrt{1-\rho^{2}}W^{2,\perp}_{t},$
where the independent standard Brownian motions $(W^{1,\perp}_{t})_{t}$ and $(W^{2,\perp}_{t})_{t}$ are such that $W$ is independent of $W^{2,\perp}$, and $W^{\beta}$ is independent of $W^{1,\perp}$. By definition, the excess return $(\beta_{t})_{t}$ reverts to its long-run average $\overline{\beta}$.
@@
We denote by $\bar{\mathbb{F}}:=(\bar{\mathcal{F}}_{t})_{t}$ the completed filtration jointly generated by $W$ and $W^{\beta}$, that is, $\bar{\mathbb{F}}:=\mathbb{F}^{W}\vee\mathbb{F}^{\beta}$, and make the following standing assumption.
@@
is an $\bar{\mathbb{F}}$- martingale under $\mathbb{P}$, with $(\mathcal{H}_{t})_{t}$ as in (2.2).
@@
    $(c_{t})_{t}$ and $(\pi_{t})_{t}$ are $\bar{\mathbb{F}}$-progressively measurable, and are such that $c_{t}\geq 0$ $\mathbb{P}$-a.s. for all $t\geq 0$, $\int_{0}^{T}c_{s}\,ds<\infty$ and $\int_{0}^{T}\pi_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$.
@@ 2.2. From a Dynamic to a Static Budget Constraint
(2.5) $\mathcal{D}:=\Big\{(D_{t})_{t\geq 0}:D\text{ is }\bar{\mathbb{F}}\text{-adapted, nonnegative, nonincreasing, càdlàg, and }D_{0^{-}}=1\;\mathbb{P}\text{-a.s.}\Big\},$
@@
    For any nonnegative $\bar{\mathbb{F}}$-progressively measurable $(c_{t})_{t}$ with $\int_{0}^{T}c_{s}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, and such that
@@
The proof borrows arguments from [24].
Proof of (1). Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$ and let $D\in\mathcal{D}$. An application of Itô’s formula for semimartingales (see Theorem 32 on p. 78 in [38]) yields
@@
Proof of (2). For $T>0$ arbitrary but fixed, consider now the reflected backward stochastic differential equation (RBSDE):
(2.11) $\begin{cases}\tilde{X}_{s}=\int_{s}^{T}\mathcal{H}_{t}(c_{t}-\ell)\,dt-\int_{s}^{T}\tilde{Z}_{t}\,dW_{t}+\int_{s}^{T}d\tilde{K}_{t},\quad\forall s\geq 0,\\ \tilde{X}_{t}\geq 0,\quad\int_{0}^{T}\tilde{X}_{t}\,d\tilde{K}_{t}=0.\end{cases}$
@@
$\mathbb{E}\left[\sup_{t\in[0,T]}|\tilde{X}_{t}|^{\nu}\right]<\infty\quad\text{and}\quad\mathbb{E}\left[\left(\int_{0}^{T}|\tilde{Z}_{t}|^{2}\,dt\right)^{\nu/2}\right]<\infty.$
Note that the driver of the RBSDE (2.11), namely $\mathcal{H}_{t}(c_{t}-\ell)$, does not depend on $\tilde{Z}$ and $\tilde{X}$. Furthermore, due to (2.10), we have for all $T>0$
$\mathbb{E}\left[\int_{0}^{T}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]<\infty.$
@@
Let then $D\in\mathcal{D}$ and define the sequence of localizing stopping times
$\vartheta_{n}:=\inf\left\{s\geq 0:\int_{0}^{s}\tilde{Z}_{u}^{2}D_{u}^{2}\,du\geq n\right\},\quad n\in\mathbb{N}.$
Note that $\vartheta_{n}\to\infty$ $\mathbb{P}$-a.s. as $n\to\infty$. Using Itô’s formula on the interval $0,T\wedge\vartheta_{n}]$ and taking expectations, we obtain
(2.12) $\tilde{X}_{0}$ $=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}D_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}D_{t-}\,d\tilde{K}_{t}-\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,dD_{t-}\right].$
Since $D$ is decreasing and $\tilde{K}$ is increasing, we have from (2.12)
$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right],$
and letting $n\to\infty$ and $T\to\infty$, (2.10) allows us to invoke the Dominated Convergence Theorem to obtain
$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right].$
Since $D\in\mathcal{D}$ was arbitrary, it follows from (2.7) that
(2.13) $\tilde{X}_{0}\geq x.$
On the other hand, let
$\tilde{D}_{s}:=\mathbf{1}_{\{s<\tilde{\eta}\}},\;s\geq 0,\quad\tilde{\eta}:=\inf\left\{s\geq 0:\tilde{X}_{s}\leq 0\right\}\wedge(T\wedge\vartheta_{n}).$
Then,
$\tilde{X}_{\tilde{\eta}}\mathbf{1}_{\{\tilde{\eta}<T\wedge\vartheta_{n}\}}=0,\quad\tilde{K}_{t}=0\quad\text{on}\quad 0\leq t<\tilde{\eta}$,
and
$\int_{0}^{T\wedge\vartheta_{n}}\tilde{D}_{t-}\,d\tilde{K}_{t}=\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,d\tilde{D}_{t-}=0,$
so that (2.12) implies
(2.14) $\tilde{X}_{0}=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\tilde{D}_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right].$
Since $\tilde{X}$ belongs to the Class D, we have $\lim_{n\to\infty}\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\right]=\mathbb{E}\left[\tilde{X}_{T}\right]=0$. Hence, using the Dominated Convergence Theorem again to take limits as $n\rightarrow\infty$ and $T\rightarrow\infty$ in (2.14), we obtain
(2.15) $\tilde{X}_{0}\leq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right]\leq x,$
where we used the fact that $\tilde{D}\in\mathcal{D}$. Therefore, combining (2.13) and (2.15), we conclude that $\tilde{X}_{0}=x$.
From the RBSDE (2.11), we know that $\tilde{X}$ satisfies
$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\tilde{Z}_{t}\,dW_{t},\quad\forall s\geq 0.$
If we now choose $\pi$ defined as
(2.16) $\pi_{t}:=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\mathcal{H}_{t}\right),\quad t\geq 0,$
@@
where $(X^{\pi,c}_{t})_{t}$ is the wealth process controlled through $(\pi,c)$, we obtain for any $s\geq 0$
@@
$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\mathcal{H}_{t}\left(\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\right)\,dW_{t}.$
We need to check that such a $\pi$ is indeed such that $(\pi,c)\in\mathcal{A}(x)$. Observe that the expression on the right-hand side of the last display equation is exactly the discounted wealth $X^{\pi,c}_{s}\mathcal{H}_{s}$. Since $\mathcal{H}_{s}$ is strictly positive for all $s\geq 0$, we thus get $X^{\pi,c}_{s}\geq 0$ for all $s\geq 0$.
Let us now check that
$\int_{0}^{T}\pi_{t}^{2}\,dt<\infty\quad\mathbb{P}-a.s.\ \text{for all}\;T>0.$
Since $X_{t}^{\pi,c}=\mathcal{H}_{t}^{-1}(\tilde{X}_{t}+\tilde{K}_{t})$, by substituting this expression into (2.16), we find

### 11 Mar 2026 v3

@@ Abstract.
In this paper, we study an intertemporal utility maximization problem in which an investor chooses consumption and portfolio strategies in the presence of a stochastic factor and a no-borrowing constraint. In the spirit of the Kim-Omberg model, the stochastic factor represents the expected excess return of the risky asset. It is perfectly negatively correlated with shocks to the risky asset, and follows an Ornstein-Uhlenbeck process, thereby capturing the mean reversion of expected excess returns–a feature well supported by empirical evidence in financial markets. The investor seeks to maximize expected utility from consumption, subject to the constraint that wealth remains nonnegative at all times. To address the dynamic no-borrowing constraint, we use Lagrange duality to transform the primal problem into a singular control problem in the dual space. We then characterize the solution to the dual singular control problem via an auxiliary two-dimensional optimal stopping problem featuring stochastic volatility, and subsequently retrieve the primal value function as well as the optimal portfolio and consumption plans. Finally, a numerical study is conducted to derive economic and financial implications.
@@ 1. Introduction
Investment opportunities vary over time because the expected excess return of the risky asset is driven by a stochastic factor $(\beta_{t})_{t}$. Empirical evidence (see, e.g., [15] and [38]) suggests that expected excess returns are predictable and mean-reverting. Motivated by this finding, we model $(\beta_{t})_{t}$ as a mean-reverting Ornstein–Uhlenbeck process, following the so-called Kim-Omberg model introduced in [27]. Furthermore, as in [41], we focus on the case in which the Brownian motion driving $(\beta_{t})_{t}$ is perfectly negatively correlated with the Brownian motion driving the risky asset’s returns, a specification that is empirically motivated by the return-predictability literature (see, e.g., [5], [15] and [38]), where variables proxying time-varying expected excess returns are found to be strongly negatively correlated with contemporaneous stock returns and positively correlated with future returns.
In such a complete market setting, we derive optimal consumption and portfolio policies, as well as regularity results for the value function. As discussed in more detail below, this is achieved by means of a Lagrange duality approach, which allows us to connect the (primal) dynamic optimization problem with a no-borrowing constraint to a singular stochastic control problem. The latter is then analyzed through an auxiliary genuinely two-dimensional optimal stopping problem, in which one of the state variables exhibits stochastic volatility. In particular, we establish continuous differentiability of the optimal stopping problem’s value function and characterize the optimal stopping time in terms of an excess-return-dependent free boundary. These properties are then instrumental in obtaining the complete solution to the original optimal consumption and portfolio choice problem.
@@
Inspired by [13], [22], and the more recent [24], we instead adopt a duality-based approach to handle the no-borrowing constraint. Instead of enforcing $X_{t}\geq 0$ dynamically at every point in time, we reformulate it as a single static budget constraint. This transformation is achieved by introducing a non-increasing, càdlàg process $(D_{t})_{t}$ that acts as an endogenous dynamic Lagrange multiplier designed to ensure $X_{t}\geq 0$ at all times $t\geq 0$ almost surely. This dual formulation allows us to express the problem in terms of an auxiliary process $(Z_{t})_{t}$ rather than the wealth process itself. Consequently, the dual problem is recast as a two-dimensional singular control problem in the state variables $(Z_{t},\beta_{t})_{t}$, where $(Z_{t})_{t}=(Z_{t}^{D})_{t}$ is controlled through the monotone control $(D_{t})_{t}$. Crucially, because the stochastic factor $(\beta_{t})_{t}$ directly drives the diffusion of the dual state, the system inherently features stochastic volatility. It is worth noticing that directly approaching the dynamic programming principle (Hamilton-Jacobi-Bellman, HJB) equation related to the primal problem is particularly challenging in our setting, because of a non-geometric wealth dynamics due to the presence of labor income. As a matter of fact, in many classical settings, the wealth process is geometric, and the positivity of wealth is automatically guaranteed. In presence of power utility functions, this geometric structure also allows the value function to be scaled with wealth, reducing the HJB equation to a one-dimensional nonlinear ODE (see, e.g., [32] and [31] for earlier works or the recent [18] and [19]). In our setting, however, the usual scaling arguments break down, the no-borrowing constraint must be enforced explicitly, and a duality approach is needed (see also Remark 2.3 below).
@@
By connecting the optimal stopping problem back to the dual singular control problem, we show that the dual value function is precisely the integral of the stopping value function with respect to the first component of the state process. This critical relationship allows us to uniquely characterize the optimal singular control $(D^{*}_{t})_{t}$, which is the minimal process (à la Skorokhod) that keeps the dual state process $(Z_{t},\beta_{t})_{t}$ within the region $\{(z,\beta):z<z^{*}(\beta)\}$. Finally, in our complete market setting, we prove that strong duality holds, demonstrating that the primal value function can be recovered from the dual value function, and vice versa. Using these results, we are then able to retrieve the optimal consumption plan $(c^{*}_{t})_{t}$, the optimal portfolio strategy $(\pi^{*}_{t})_{t}$, and the optimal wealth process $(X^{*}_{t})_{t}$, which, as required, satisfies the dynamic no-borrowing constraint.
@@
Optimal consumption and investment problems in the presence of a stochastic factor have also been widely studied. For example, [33] consider an optimal consumption problem involving stochastic interest rates, while [20] study a Merton consumption and portfolio problem with stochastic asset returns and volatilities (see also [21]). Closely related to our work is the Kim-Omberg model [27], which studies the dynamic non-myopic portfolio behavior of an investor trading a risk-free and a risky asset, with expected excess returns following a mean-reverting Ornstein-Uhlenbeck process. Extending this framework to include intermediate consumption, [41] appears to be the first to derive exact optimal portfolio and consumption rules in this setting. However, unlike our model, [41] does not incorporate labor income, works with a finite-time horizon and enforces the non-negativity constraint on wealth only at the terminal time $T$. Furthermore, [18] derive optimal consumption and investment policies in a complete market featuring a stochastic factor, modeled as a general scalar diffusion that drives investment opportunities. In their framework, the absence of labor income ensures that the wealth dynamics remain purely geometric. This in turn allows the value function to be scaled by wealth, thereby reducing the dimension of the HJB equation. Lastly, [19] consider an optimal consumption and investment model with general stochastic factors. They address optimal investment and consumption for a power utility investor using an incomplete stochastic factor model on an infinite horizon and provide a complete characterization for a finite state space. When the factor follows a diffusion process, they develop a new theoretical framework to prove existence and bound the HJB solution, verifying models like the Heston model rigorously for the first time. Conversely to our approach, [19] allows for general stochastic factors, including the Kim-Omberg setting, but assumes geometric wealth without labor income, which automatically enforces the non-negativity of wealth and crucially permits dimension reduction of the associated HJB equation.
@@ 2.1. The Financial Market and the Agent’s Problem
(2.1) $dS_{t}=(r+\beta_{t})S_{t}\,dt+\sigma S_{t}\,dW_{t},\quad t>0,\quad S_{0}=s>0,$
@@
(2.2) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
where $\kappa>0$ is the speed of mean reversion, $\overline{\beta}>0$ is the equilibrium level, and $\sigma_{\beta}>0$ is the volatility of $(\beta_{t})_{t}$. By definition, the expected excess return $(\beta_{t})_{t}$ reverts to its long-run average $\overline{\beta}$. As one can see from (2.1) and (2.2), $(S_{t})_{t}$ and $(\beta_{t})_{t}$ are perfectly negatively correlated. This complete-market specification is in the spirit of [41], where perfect negative correlation is also assumed in a Kim-Omberg setting, and it plays a fundamental role in the subsequent mathematical analysis. Empirically, this assumption is motivated by the evidence that variables proxying time-varying expected excess returns, such as the dividend-price ratio, are strongly negatively correlated with contemporaneous stock returns and positively correlated with future returns (cf. [5], [15] and [38]).
@@
We make the following standing assumption.
@@
is an $\mathbb{F}^{W}$- martingale under $\mathbb{P}$, with $(\mathcal{H}_{t})_{t}$ as in (2.4).
@@
    $(c_{t})_{t}$ and $(\pi_{t})_{t}$ are $\mathbb{F}^{W}$-progressively measurable, and are such that $c_{t}\geq 0$ $\mathbb{P}$-a.s. for all $t\geq 0$, $\int_{0}^{T}c_{s}\,ds<\infty$ and $\int_{0}^{T}\pi_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$.
@@ 2.2. From a Dynamic to a Static Budget Constraint
(2.7) $\mathcal{D}:=\Big\{(D_{t})_{t\geq 0}:D\text{ is }\mathbb{F}^{W}\text{-adapted, nonnegative, nonincreasing, càdlàg, and }D_{0^{-}}=1\;\mathbb{P}\text{-a.s.}\Big\},$
@@
    For any nonnegative $\mathbb{F}^{W}$-progressively measurable $(c_{t})_{t}$ with $\int_{0}^{T}c_{s}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, and such that
@@
Proof of (1). The proof borrows arguments from [24].
Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$ and let $D\in\mathcal{D}$. An application of Itô’s formula for semimartingales (see Theorem 32 on p. 78 in [39]) yields
@@
Proof of (2). We set
(2.13) $A_{t}:=\int_{0}^{t}\mathcal{H}_{s}(c_{s}-\ell)ds,\quad t\geq 0,$
and have for any $\mathbb{F}^{W}$-stopping time $\eta$
$|A_{\eta}|\leq\int_{0}^{\infty}\mathcal{H}_{s}|c_{s}-\ell|ds.$
From (2.12), which holds thanks to (2.9) and can be shown as in the proof of (1), it then follows that the process $(A_{t})_{t}$ is continuous and belongs to Class D [^1] A process $(\zeta_{t})_{t}$ is of Class (D) if the family of random variables $\left\{\zeta_{\tau}:\tau<\infty\;\mathbb{P}\text{-a.s.\ is a stopping time}\right\}$ is uniformly integrable..
Now fix an almost surely finite $\mathbb{F}^{W}$-stopping time $\eta$ and define
$D_{t}^{\eta}:=\mathbf{1}_{\{t<\eta\}},\qquad t\geq 0,$
so that $D^{\eta}\in\mathcal{D}$, which implies by (2.9)
$\mathbb{E}\left[\int_{0}^{\eta}\mathcal{H}_{t}(c_{t}-\ell)\,dt\right]\leq x,$
or equivalently
(2.14) $\mathbb{E}[A_{\eta}]\leq x,$
upon using (2.13). We then define the Snell envelope (cf., e.g., page 8 in [36]) of $(A_{t})_{t}$ by
$Y_{t}:=\operatorname*{ess\,sup}_{\eta\geq t}\mathbb{E}[A_{\eta}\mid\mathcal{F}^{W}_{t}],\qquad t\geq 0.$
Since $(A_{t})_{t}$ is continuous and belongs to Class D, it thus follows that $(Y_{t})_{t\geq 0}$ is a continuous supermartingale belonging to Class D, and it is the smallest supermartingale dominating $(A_{t})_{t}$; that is,
(2.15) $Y_{t}\geq A_{t},\qquad\text{for all }t\geq 0.$
Hence, (2.14) implies
(2.16) $Y_{0}=\sup_{\eta\geq 0}\mathbb{E}[A_{\eta}]\leq x.$
Next, we define the continuous supermartingale
$\bar{Y}_{t}:=Y_{t}+(x-Y_{0}),\qquad t\geq 0,$
which, thanks to (2.15) and (2.16), satisfies
$\bar{Y}_{0}=x\qquad\text{and}\qquad\bar{Y}_{t}\geq A_{t}\quad\text{for all }t\geq 0.$
By the Doob-Meyer decomposition (cf., e.g., Theorem 3.1 on page 56 in [36]), there exists a uniformly integrable continuous martingale $(\bar{N}_{t})_{t\geq 0}$ and a continuous predictable increasing process $(K_{t})_{t\geq 0}$ with $K_{0}=0$ such that
(2.17) $\bar{Y}_{t}=x+\bar{N}_{t}-K_{t},\qquad t\geq 0.$
Moreover, the martingale representation theorem (cf., e.g., Theorem 4.3.4 on page 53 in [35]) yields a progressively measurable process $(\tilde{Z}_{t})_{t\geq 0}$ such that
(2.18) $\bar{N}_{t}=\int_{0}^{t}\tilde{Z}_{s}\,dW_{s},\qquad t\geq 0,$
and
@@
We now define
(2.19) $\widehat{X}_{t}:=\bar{Y}_{t}-A_{t}+K_{t},\qquad t\geq 0.$
Since $\bar{Y}_{t}\geq A_{t}$ and $K_{t}\geq 0$, we have
$\widehat{X}_{t}\geq 0,\qquad t\geq 0,$
and, upon using $K_{0}=A_{0}=0$ and $\bar{Y}_{0}=x$, also $\widehat{X}_{0}=x$. Plugging (2.13) and (2.17) into (2.19) yields
$\widehat{X}_{t}=x+\bar{N}_{t}-K_{t}-\int_{0}^{t}\mathcal{H}_{s}(c_{s}-\ell)\,ds+K_{t}=x+\int_{0}^{t}\mathcal{H}_{s}(\ell-c_{s})\,ds+\int_{0}^{t}\tilde{Z}_{s}\,dW_{s},$
where we have used (2.18) in the final display equation above, as well. Therefore,
(2.20) $d\widehat{X}_{t}=\mathcal{H}_{t}(\ell-c_{t})\,dt+\tilde{Z}_{t}\,dW_{t},\qquad\widehat{X}_{0}=x.$
Let us now construct the candidate investment policy $(\pi_{t})_{t}$ as follows
(2.21) $\pi_{t}:=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\widehat{X}_{t}\right),\qquad t\geq 0,$
which is clearly $\mathbb{F}^{W}$-progressively measurable.
In order to show
@@
we note that the processes $(\mathcal{H}_{t})_{t}$, $(\mathcal{H}^{-1}_{t})_{t}$, $(\beta_{t})_{t}$, and $(\widehat{X}_{t})_{t}$ are continuous, and hence pathwise bounded. Since we also have $\int_{0}^{T}\tilde{Z}_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, we obtain (2.22) from (2.21).
@@
It remains to verify that $(X^{\pi,c}_{t})_{t\geq 0}$ is such that $X^{\pi,c}_{t}\geq 0$ for all $t\geq 0$, where $(X^{\pi,c}_{t})_{t\geq 0}$ denotes the wealth controlled through $(\pi,c)$. As the process $(\mathcal{H}_{t})_{t}$ (cf. (2.4)) is such that
$d(\mathcal{H}_{t}^{-1})=\mathcal{H}_{t}^{-1}\left(r+\frac{\beta_{t}^{2}}{\sigma^{2}}\right)dt+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\,dW_{t},$
applying Itô’s formula to the process $(\mathcal{H}_{t}^{-1}\widehat{X}_{t})_{t}$ yields
$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})$ $=\mathcal{H}_{t}^{-1}d\widehat{X}_{t}+\widehat{X}_{t}\,d(\mathcal{H}_{t}^{-1})+d[\widehat{X},\mathcal{H}^{-1}]_{t}$
@@
$=(\ell-c_{t})\,dt+\mathcal{H}_{t}^{-1}\tilde{Z}_{t}\,dW_{t}+\mathcal{H}_{t}^{-1}\widehat{X}_{t}\left(r+\frac{\beta_{t}^{2}}{\sigma^{2}}\right)dt+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\,dW_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\tilde{Z}_{t}\,dt,$
where (2.20) has also been used. Rearranging terms gives